solve the given equations using exact differential

equations.

6. (cos x ∙ cos y − cot x)dx − (sin x ∙ sin y)dy = 0


7. x(3xy − 4y3 + 6)dx + (x3 − 6x2y2 − 1)dy = 0, when x = 2, y = 0


8. 2xydx + (y2 + x2)dy = 0


9. (xy2 + y − x)dx + x(xy + 1)dy = 0, when x = 1, y = 1


|10. (1 − xy)−2dx + [y2 + x2(1 − xy)−2]dy = 0, when x = 3, y = 4


1
Expert's answer
2021-09-24T06:38:02-0400

"\displaystyle
(\cos x. \cos y- \cot x)dx - (\sin x \sin y)dy= 0\\
\text{Let $M = \cos x. \cos y- \cot x$ and $N = -\sin x \sin y$}\\
\text{Let $F(x,y) = \int (-\sin x \sin y)dy$}\\
\implies F(x,y) = \sin x \cos y + h(y)\\
\text{Differentiating and equating it to M, we have}\\
\cos x \cos y +h'(x) = \cos x \cos y -\cot x\\
\implies h'(x) = -\cot x\\
h(x) = -\ln \sin x\\
\therefore F(x,y) = \sin x \cos y -\ln \sin x\\
\text{The solution to the exact differential equation is}\\
\sin x \cos y -\ln \sin x=c\\
2. 
(3x^2y-4y^3+6)dx - (x^3-6x^2y^2-1)dy= 0\\
\text{Let $M = 3x^2y-4y^3+6$ and $N = x^3-6x^2y^2-1$}\\
\text{Let $F(x,y) = \int (x^3-6x^2y^2-1)dy$}\\
\implies F(x,y) = x^3y-2x^2y^3-y + h(y)\\
\text{Differentiating and equating it to M, we have}\\
3x^2y-4y^3+h'(x) = 3x^2y-4y^3+6\\
\implies h'(x) = 6\\
h(x) = 6x\\
\therefore F(x,y) = x^3y-2x^2y^3-y +6x\\
\text{The solution to the exact differential equation is}\\
x^3y-2x^2y^3-y +6x=c\\
\text{When x = 2, y=0, hence c = 12}\\
\implies x^3y-2x^2y^3-y +6x=12
3.
(2xy)dx - (y^2+x^2)dy= 0\\
\text{Let $M = 2xy$ and $N = y^2+x^2$}\\
\text{Let $F(x,y) = \int (y^2+x^2)dy$}\\
\implies F(x,y) = \frac{y^3}{3}+x^2y + h(x)\\
\text{Differentiating and equating it to M, we have}\\
2xy +h'(x) = 2xy\\
\implies h'(x) = 0\\
h(x) = c\\
\therefore F(x,y) = \frac{y^3}{3}+x^2y+c\\
\text{The solution to the exact differential equation is}\\
\frac{y^3}{3}+x^2y+c=c\\
\implies \frac{y^3}{3}+x^2y+c=0
4.
(xy^2+y-x)dx - (x^2y+x)dy= 0\\
\text{Let $M = xy^2+y-x$ and $N = x^2y+x$}\\
\text{Let $F(x,y) = \int (x^2y+x)dy$}\\
\implies F(x,y) = \frac{x^2y^2}{2}+xy + h(x)\\
\text{Differentiating and equating it to M, we have}\\
xy^2+y+h'(x) = xy^2+y-x\\
\implies h'(x) = -x\\
h(x) = -\frac{x^2}{2}\\
\therefore F(x,y) = x^3y-2x^2y^3-y +6x\\
\text{The solution to the exact differential equation is}\\
\frac{x^2y^2}{2}+xy-\frac{x^2}{2}=c\\
\text{When x = 1, y=1, hence c = 1}\\
\implies \frac{x^2y^2}{2}+xy-\frac{x^2}{2}=1\\
5. 
(1-xy-2)dx + (y^2+x^2-x^3y-2)dy\\
\text{The equation isn't exact as}\\
M_y =-x, N_x=2x-3x^2y\\
M_y \neq N_x"